In 1999, the US produced about 2 million more motor vehicles than Japan. Together, the two countries produced about 22 million motor vehicles. how may vehicles were produced in each country?
Answer by checkley75(3666)

X+(X+2,000,000)=22,000,000
2X=22,000,000-2,000,000
2X=20,000,000
X=20,000,000/2
X=10,000,000 NUMBER OF CARS PRODUCED IN JAPAN
22,000,000-10,000,000=12,000,000 NUMBER OF CARD PRODUCED IN US.